At 25 °C, only 0.0140 mol of the generic salt AB3 is soluble in 1.00 L of water.

What is the
1
Expert's answer
2020-04-27T01:40:43-0400

Correct task : At 25°C, only 0.0140 mol of the generic salt AB3 is soluble in 1.00 L of water. What is the Ksp of this salt at 25°C.


Solution:

When the salt dissolves, it dissociates as follows:

AB3 --> A3+ + 3B⁻

--S--------S-------3S--


The molar solubility (S) is the number of moles that can dissolve in 1 L of solution.

Molar solubility of salt is (0.0140 mol) / (1.00 L) = 0.0140 mol/L.

According to the dissociation equation:

Solubility of A3+ is 0.0140 mol/L and solubility of B⁻ is 3×0.0140 mol= 0.0420 mol/L.

[A3+] = 0.0140 mol/L.

[B⁻] = 0.0420 mol/L.


Ksp is the solubility product constant and calculated as follows:

Ksp(AB3) = [A3+] × [B⁻]3

Ksp(AB3) = [0.0140] × [0.0420]3

Ksp(AB3) = 10.37×10-7.

Ksp of this salt is 1.04×10-6.


Answer: 1.04×10-6 is the Ksp of this salt at 25°C.